# Pro-Prophet: Systematic Load Balancing Method for Efficient Parallel Training of Large-scale MoE Models

### v1

@@ Abstract
The size of deep learning models has been increasing to enhance model quality. The linear increase in training computation budget with model size means that training an extremely large-scale model is exceedingly time-consuming. Recently, the Mixture of Expert (MoE) has drawn significant attention as it can scale models to extra-large sizes with a stable computation budget. However, inefficient distributed training of large-scale MoE models hinders their broader application. Specifically, a considerable dynamic load imbalance occurs among devices during training, significantly reducing throughput. Several load-balancing works have been proposed to address the challenge. System-level solutions draw more attention for their hardware affinity and non-disruption of model convergence compared to algorithm-level ones. However, they are troubled by high communication costs and poor communication-computation overlapping. To address these challenges, we propose a systematic load-balancing method, Pro-Prophet, which consists of a planner and a scheduler for efficient parallel training of large-scale MoE models. To adapt to the dynamic load imbalance, we profile training statistics and use them to design Pro-Prophet. For lower communication volume, Pro-Prophet planner determines a series of lightweight load-balancing strategies and efficiently searches for a communication-efficient one for training based on the statistics. For sufficient overlapping of communication and computation, Pro-Prophet scheduler schedules the data-dependent operations based on the statistics and operation features, further improving the training throughput. We conduct extensive experiments in four clusters and five MoE models. The results indicate that Pro-Prophet achieves up to 2.66x speedup compared to two popular MoE frameworks including Deepspeed-MoE and FasterMoE. Additionally, Pro-Prophet achieves a load-balancing enhancement of up to 11.01x when compared to a representative load-balancing work, FasterMoE.
@@ I Introduction
However, these systematic solutions struggle to enhance training efficiency effectively due to two drawbacks. 1) Heavy communications of model states (i.e., parameters, gradients and optimizer states [10]) are introduced. The expert placement proposed by previous works introduces a global transfer of parameters and gradients or a whole model states communication. These strategies involve unnecessary communications across devices, hindering the improvement of training efficiency. 2) Devices experience significant communication and computation idle during training. Due to data dependencies among operators, the solutions have to perform some communications and computations sequentially. For example, only the experts have been selected and their model states have been transmitted, their computations of inputs can be launched. And then, the aggregation of gradients occurs only after the computation of gradients is finished. They neglect the potential of communication and computation overlapping, thus significantly influencing device utilization.
In this paper, we present a systematic load-balancing solution, Pro-Prophet, which overcomes two drawbacks by a planner and scheduler respectively.
@@
In contrast to these systematic works, Pro-Prophet planner introduces a series of lightweight expert placements. In a lightweight expert placement, each expert is independently allocated to a subset of devices. Communication of parameters and gradients for the expert occurs in these specific devices. Serve to select an optimal expert placement, the planner proposes a performance model that estimates the execution time of a MoE layer employing a lightweight expert placement. However, it is non-trivial to find it due to the combinatorial explosion of the number of expert placements. To tackle this, the planner designs a locality-based greedy algorithm. The algorithm employs a greedy strategy to search for a communication-efficient expert placement. Besides, its launching frequency is reduced based on the locality, further improving the training throughput.
@@ II-A Background
In recent years, dynamic sparse-activated architectures have been proposed to solve the trouble. One of the popular approaches is the Mixture of Experts (MoE), which can significantly improve the model capacity while maintaining a consistent computational budget. Nowadays, MoE has been successfully applied to large pre-training language models. Excellent MoE models that appeared in industry and academia greatly draw researchers’ attention. For example, Google has trained a series of MoE models called Glam[2]. The largest Glam model is seven times larger than GPT-3, but the training cost is less than 1/3 of it. Experiments show that these models achieve higher accuracy than GPT-3 in 29 zero, single, and small sample learning tasks, representing the superiority of MoE models. The other example is GPT-4[19]. The technical report of OpenAI indicates that the GPT-4 is a MoE model, which achieved the highest performance in various downstream tasks. Besides, the ChatGPT based on the GPT-4 has caused a tremendous sensation.
@@
Fig. 2 illustrates a workflow of EP in a MoE layer. Firstly, the gate network determines top-$1$ expert for each input. Then inputs are transferred to corresponding devices via an All-to-All (A2A) communication operation. Subsequently, each device performs the expert computation for collected inputs and then launches another A2A to reorganize the results back to the inputs’ original devices for computations or communications of the subsequent non-MoE layer. Nowadays, many popular distributed frameworks support the training of large-scale MoE models using EP[20, 4, 21, 22, 23].
@@ II-B Motivation
TABLE I: Time breakdown of training. Load balancing is short for L.b.
@@
As shown in Table I, previous solutions introduce a Search, Place and Reduce processes to balance the load. However, the overhead of load-balancing is up to 37.1%. There are two reasons behind this huge cost. Firstly, they do heavy communication of model states. They have to transfer the parameters and gradients of heavy-load experts among all devices or transmit the whole model states of experts. Secondly, they cannot sufficiently overlap the communication and computations during the training of a MoE model. Due to the data-dependency, they have to do communication and computation sequentially.
@@ III Overview of Pro-Prophet
Motivated by Section II, we propose a systematical load-balancing approach, Pro-Prophet, which can efficiently balance the load of devices. The overview of Pro-Prophet is presented in Fig. 5. Pro-Prophet is composed of a planner and a scheduler. MoE model, locality and device pool are three inputs of it. The device pool defines the topology of devices. The utilization of the locality is the key advantage of Pro-Prophet.
@@
Figure 5: The overview of Pro-Prophet. Pro-Prophet is composed of Pro-Prophet planner and Pro-Prophet scheduler. MoE model, locality and device pool are three inputs of it. Firstly, Pro-Prophet planner searches for a communication-efficient expert placement using its locality-based greedy algorithm. The algorithm iteratively generates and evaluates a lightweight expert placement utilizing its performance model until the load is balanced. Then the execution engine produces a load-balancing workflow based on the planner. Finally, Pro-Prophet scheduler schedules three data-dependent operations to parallel operations for communication and computation overlapping, further improving the training throughput.
@@ IV-A Lightweight Expert Placement
In a lightweight expert placement, each expert is mapped to one or more devices independently. Only the parameters and gradients rather than all model states are transferred among its devices. We use Trans and Agg primitives to describe these two communications respectively. In the forward pass, a Trans is first launched to transfer the parameters. After that, each device contains the parameters of some experts, thus its local inputs routed to these experts could be computed locally. After the backward computation, the gradients of an expert could be generated in several devices. As each device only maintains the optimizer states of one expert, a Agg primitive is launched to aggregate gradients of each expert to its original device. This design has two advantages: 1) Only part of the model states are communicated. 2) The model states are only communicated among a subset of devices.
@@ 24       $n\_bottom\leftarrow BottomK(gating,n)$ ;
28      $n\_bottoms.append(n_{b}ottom)$
@@ 29      $s\leftarrow size(L)$
30       $H,R\leftarrow Replace\_Inputs(L,n_{b}ottoms)$
@@ V-B Block-wise scheduling strategy
A dynamic scheduling strategy is beneficial as the load of devices fluctuates with the training process. However, the non-negligible overhead will be introduced if we determine an optimal scheduling strategy that as far as possible hides the overhead of three types of primitives involved by systematical load balancing methods in the runtime. Therefore, we design an offline scheduling policy to overlap communication and computation while avoiding the extra overhead. The policy is founded on static elements within dynamics, not intuition. Our block-wise scheduling strategy is summarized in Algorithm 2. The first primitive is Plan. We schedule the Plan primitive of the $i$-th block in the iteration $j+1$ to the A2A communication of the $i$-th block in the $j$-th iteration. For the Trans primitive, we overlap it with computations of the former block within an iteration. Specifically, the forward computation of the $i$-th block is responsible for overlapping the Trans primitive of the $i+1$-th block. As two computations are executed in the $i$-th block, we split the Trans primitive into two sub-primitives and launch them simultaneously with the two computations. Even though the duration of Trans and EFC varies as the device loads, the forward computation overhead of the non-MoE layer and the transferring overhead of an expert’s parameters are static. We can estimate them before training and properly split the Trans primitive. The advantage of the estimation is that we can exhaustively fill in the communication idle in the performing of the forward computation of the non-MoE layer. Finally, we schedule the $Agg$ of block $i+1$ into the backward computation of the $i$-th one. Similarly, we can estimate the backward computation overhead of the non-MoE layer and do a suitable communication partition.
@@ VII Related Work
Communication schedule. Overlapping communications and computing can enhance hardware utilization and improve the system’s throughput [46, 47]. Previous communication scheduling methods [27, 48, 5] for dense models have demonstrated promising results. In this paragraph, We focus on introducing works designed for MoE models.

### v2

# Pro-Prophet: A Systematic Load Balancing Method for Efficient Parallel Training of Large-scale MoE Models

@@ Abstract
The size of deep learning models has been increasing to enhance model quality. The linear increase in training computation budget with model size means that training an extremely large-scale model is exceedingly time-consuming. Recently, the Mixture of Expert (MoE) has drawn significant attention as it can scale models to extra-large sizes with a stable computation budget. However, inefficient distributed training of large-scale MoE models hinders their broader application. Specifically, a considerable dynamic load imbalance occurs among devices during training, significantly reducing throughput. Several load-balancing works have been proposed to address the challenge. System-level solutions draw more attention for their hardware affinity and non-disruption of model convergence compared to algorithm-level ones. However, they are troubled by high communication costs and poor communication-computation overlapping. To address these challenges, we propose a systematic load-balancing method, Pro-Prophet, which consists of a planner and a scheduler for efficient parallel training of large-scale MoE models. To adapt to the dynamic load imbalance, we profile training statistics and use them to design Pro-Prophet. For lower communication volume, Pro-Prophet planner determines a series of lightweight load-balancing strategies and efficiently searches for a communication-efficient one for training based on the statistics. For sufficient overlapping of communication and computation, Pro-Prophet scheduler schedules the data-dependent operations based on the statistics and operation features, further improving the training throughput. We conduct extensive experiments in four clusters and five MoE models. The results indicate that Pro-Prophet achieves up to 2.66x speedup compared to two popular MoE frameworks including Deepspeed-MoE and FasterMoE. Furthermore, Pro-Prophet has demonstrated a load-balancing improvement of up to 11.01x compared to a representative load-balancing work, FasterMoE.
@@ I Introduction
However, these systematic solutions struggle to enhance training efficiency effectively due to two drawbacks. 1) Heavy communications of model states (i.e., parameters, gradients and optimizer states [10]) are introduced. The previous expert placements introduce a global transfer of parameters and gradients or a whole model states communication. These transferring strategies involve unnecessary communications across devices, hindering the improvement of training efficiency. 2) Devices experience significant communication and computation idle during training. Due to data dependencies among operators, the solutions have to perform some communications and computations sequentially. For example, only the experts have been selected and their model states have been transmitted, their computations of inputs can be launched. And then, the aggregation of gradients occurs only after the computation of gradients is finished. They neglect the potential of communication and computation overlapping, thus significantly influencing device utilization.
In this paper, we propose a systematic load-balancing solution, Pro-Prophet, which overcomes two drawbacks by a planner and scheduler respectively.
@@
To reduce the communication volume, Pro-Prophet planner introduces a series of lightweight expert placements. In a lightweight expert placement, each expert is independently allocated to a subset of devices. Communication of parameters and gradients for the expert occurs in these specific devices. Serve to evaluate expert placements, the planner proposes a performance model that estimates the execution time of a MoE layer employing a lightweight expert placement. However, it is non-trivial to find the optimal one due to the combinatorial explosion of the number of expert placements. To tackle this, the planner designs a locality-based greedy algorithm. The algorithm employs a greedy strategy to search for a communication-efficient expert placement. Besides, its launching frequency is reduced based on the locality, further improving the training throughput.
@@ II-A Background
In recent years, dynamic sparse-activated architectures have been proposed to solve the trouble. One of the popular approaches is the Mixture of Experts (MoE), which can significantly improve the model capacity while maintaining a consistent computational budget. Nowadays, MoE has been successfully applied to large language models[19, 20, 21, 22, 23, 24, 25]. Excellent MoE models that appeared in industry and academia greatly draw researchers’ attention. For example, Google has trained a series of MoE models called Glam[2]. The largest Glam model is seven times larger than GPT-3, but the training cost is less than 1/3 of it. Experiments show that these models achieve higher accuracy than GPT-3 in 29 zero, single, and small sample learning tasks, representing the superiority of MoE models. The other example is GPT-4[26]. The technical report of OpenAI indicates that the GPT-4 is a MoE model, which achieved the highest performance in various downstream tasks. Besides, the ChatGPT based on the GPT-4 has caused a tremendous sensation.
@@
Fig. 2 illustrates a workflow of EP in a MoE layer. Firstly, the gate network determines top-$1$ expert for each input. Then inputs are transferred to corresponding devices via an All-to-All (A2A) communication operation[27, 28, 29]. Subsequently, each device performs the expert computation for collected inputs and then launches another A2A to reorganize the results back to the inputs’ original devices for computations of the subsequent non-MoE layer. Nowadays, many popular distributed frameworks support the training of large-scale MoE models using EP[30, 4, 31, 32, 33].
@@ II-B Motivation
TABLE I: Time breakdown of training. L.B. is short for Load balancing
@@
As shown in Table I, previous solutions introduce a Search, Place and Reduce processes to balance the load. However, the overhead of load-balancing is up to 37.1%. There are two reasons behind this huge cost. Firstly, they do heavy communication of model states. They have to transfer the parameters and gradients of heavy-load experts among all devices or transmit the whole model states of experts. Secondly, they cannot sufficiently overlap the communication and computations during the training of a MoE model. Due to the data dependency, they have to do communication and computation sequentially.
@@ III Overview of Pro-Prophet
Motivated by Section II, we propose a systematical load-balancing approach, Pro-Prophet, which can efficiently balance the load of devices. The overview of Pro-Prophet is presented in Fig. 5. Pro-Prophet is composed of a planner and a scheduler. MoE model, locality, and device pool are three inputs of it. The device pool defines the topology of devices. The utilization of the locality is the key advantage of Pro-Prophet.
@@
Figure 5: The overview of Pro-Prophet. Pro-Prophet is composed of Pro-Prophet planner and Pro-Prophet scheduler. MoE model, locality, and device pool are three inputs of it. Firstly, Pro-Prophet planner searches for a communication-efficient expert placement using its locality-based greedy algorithm. The algorithm iteratively generates and evaluates a lightweight expert placement utilizing its performance model until the load is balanced. Then the execution engine produces a load-balancing workflow based on the planner. Finally, Pro-Prophet scheduler schedules three data-dependent operations to parallel operations for communication and computation overlapping, further improving the training throughput.
@@ IV-A Lightweight Expert Placement
In a lightweight expert placement, each expert is mapped to one or more devices independently. Only the parameters and gradients rather than all model states are transferred among its devices. We use Trans and Agg primitives to describe these two communications respectively. In the forward pass, a Trans is first launched to transfer the parameters. After that, each device contains the parameters of some expert, thus its local inputs routed to these experts could be computed locally. After the backward computation, the gradients of an expert could be generated in several devices. As each device only maintains the optimizer states of one expert, a Agg primitive is launched to aggregate gradients of each expert to its original device. This design has two advantages: 1) Only part of the model states are communicated. 2) The model states are only communicated among a subset of devices.
@@ 24       $n\_bottom\leftarrow BottomK(gating,n)$ ;
28      $n\_bottoms.append(n\_bottom)$
@@ 29      $s\leftarrow size(L)$
30       $H,R\leftarrow Replace\_Inputs(L,n\_bottoms)$
@@ V-B Block-wise scheduling strategy
A dynamic scheduling strategy is beneficial as the load of devices fluctuates with the training process. However, the non-negligible overhead will be introduced if we determine an optimal scheduling strategy that as far as possible hides the overhead of three types of primitives involved by systematical load balancing methods in the runtime. Therefore, we design an offline scheduling policy to overlap communication and computation while avoiding the extra overhead. The policy is founded on static elements within dynamics, not intuition. Our block-wise scheduling strategy is summarized in Algorithm 2. The first primitive is Plan. We schedule the Plan primitive of the $i$-th block in the iteration $j+1$ to the A2A communication of the $i$-th block in the $j$-th iteration. For the Trans primitive, we overlap it with computations of the former block within an iteration. Specifically, the forward computation of the $i$-th block is responsible for overlapping the Trans primitive of block $i+1$. As two computations are executed in the $i$-th block, we split the Trans primitive into two sub-primitives and launch them simultaneously with the two computations. Even though the duration of Trans and EFC varies as the device loads, the forward computation overhead of the non-MoE layer and the transferring overhead of an expert’s parameters are static. We can estimate them before training and properly split the Trans primitive. The advantage of the estimation is that we can exhaustively fill in the communication idle in the performing of the forward computation of the non-MoE layer. Finally, we schedule the $Agg$ of block $i+1$ into the backward computation of the $i$-th one. Similarly, we can estimate the backward computation overhead of the non-MoE layer and do a suitable communication partition.
@@ VII Related Work
Communication schedule. Overlapping communications and computing can enhance hardware utilization and improve the system’s throughput [56, 57]. Previous communication scheduling methods [37, 58] for dense models have demonstrated promising results. In this paragraph, We focus on introducing works designed for MoE models.